Can someone help me with this problem.
The number of tickets sold each day for an upcoming performance of Handel's Messiah is given by n(x) = -0.5x^2+13x+11, where x is the number of days since the concert was first announced. When will daily tickets sales peak and how many tickets will be sold that day.

Answer by stanbon(75887)

n(x) = -0.5x^2+13x+11
-----
Max occurs when x = -b/(2a) = -13/(2*-0.5) = 13/1 = 13
-------
Maximum number of tickets = n(13) = -(1/2)13^2+13*13 + 11
= (-1/2)*169 + 169 + 11
-------
= 96 when rounded up to the nearest whole number
================================================
Maximum sales occurs on day 13.
# of tickets sold that day is 96 if rounded up
# of tickets sold that day is 95 if rounded down
Cheers,
Stan H.
=============